What is the mass, in ?Sun, , of a black hole whose Schwarzschild radius is 6.8×102 km? (Round your answer to two significant figures. For example, 182.4 would be rounded to 180.)
Answer #1

Given : Rs = 6.8 ×102 km =6.8×105 m

Constant : G = 6.67×10-11 m3kg-1s-2 ; c = 3×108 m/s

Solution :

The Schwarzschild radius (Rs) is related to Msun by the following equation:

Rearranging to find mass of sun:

Putting values to get :

= 4.59×1032 kg

Answer : 4.59×1032 kg ( or 5.00 ×1032 kg Rounded)